Find the amount of money in an account after 8 years if $4800 is deposited at 6% annual interest

Answer by stanbon(75887)

Find the amount of money in an account after 8 years if $4800 is deposited at 6% annual interest
----------------------------
A(t) = P + Prt
A(t) = P(1+rt)
----
A(8) = 4800(1+0.06*8)
A(8) = 4800(1.48)
A(8) = $7104.00
======================
Cheers,
Stan H.